# Application of the FRADO model of BLR formation to the Seyfert galaxy NGC 5548 and the first step toward determining the Hubble constant

### v2

@@ Abstract
Methods. We adopt the standard accretion disk model, with the option of an inner hot flow, and employ the lamp-post model to account for disk irradiation. Additionally, we model the BLR structure based on radiation pressure acting on dust. The model is parameterized by the black hole mass (which is fixed), the accretion rate, the viewing angle, the height of the lamp-post, the cloud density, and the cloud covering factor. The resulting continuum time delays arise from a combination of disk reprocessing and the reprocessing of a fraction of radiation by the BLR.
Results. Our model reasonably reproduces the observed broad-band continuum, the H time delay, and the continuum inter-band time delays measured during the observational campaign. When the accretion rate is not constrained by the known distance to the source, our approach allows for a direct estimation of the distance. The resulting Hubble constant, = km s-1 Mpc-1, represents a significant improvement over the findings of Cackett et al. (2007).
@@
Seyfert 1 galaxy NGC 5548 is one of the most extensively observed active galactic nuclei (AGNs). Its brightness, relatively close distance, and significant variability made it an early candidate for studying the inner, unresolved structure of an active nucleus (e.g., Clavel et al., 1991; Peterson et al., 1991; Krolik et al., 1991; Rokaki et al., 1993). Among the various wavelengths, X-ray variability is the fastest, likely originating from the most compact region near the black hole. The reprocessing of this X-ray emission at longer wavelengths, observed through light echo studies, provides insights into the structure and dynamics of the surrounding material, a technique known as reverberation mapping (RM) (Blandford & McKee, 1982; Peterson, 1993). RM has been widely applied to AGNs (Sergeev et al., 2005; McHardy et al., 2014; Mudd et al., 2018; Homayouni et al., 2019; Yu et al., 2020; Guo et al., 2022), particularly in studies of continuum variability, which is believed to arise from the irradiated accretion disk. The simple standard disk model of Shakura & Sunyaev (1973) predicts that the time delays () should scale with wavelength () as . While observational results have roughly confirmed this trend, they have not matched the predicted normalization, as observed delays continue to exceed theoretical expectations (e.g. Cackett et al., 2007; Shappee et al., 2014; Kokubo, 2018; Guo et al., 2022). This discrepancy, known as the accretion disk size problem, was first recognized in optical microlensing studies of gravitationally lensed quasars (Morgan et al., 2010) and remains an ongoing challenge in our understanding of AGN accretion disk structure.
@@
The central components of AGNs include a massive black hole, a compact X-ray emitting region, a relatively cold Keplerian accretion disk, and the Broad Line Region (BLR) (see, e.g., Krolik, 1999, for a basic compendium). The Keplerian disk serves as the primary source of continuum emission, while the BLR predominantly produces broad emission lines, along with a minor contribution to the continuum. In terms of spatial extent, the innermost regions of the accretion disk span approximately 10100 , whereas the BLR extends outward to nearly 1000 .
@@
Recent intensive RM campaigns (Edelson et al., 2015; De Rosa et al., 2015; Fausnaugh et al., 2016; Pei et al., 2017; Horne et al., 2021; Lu et al., 2022) have significantly advanced our understanding of AGN nuclear structure. These studies have revealed that the observed optical continuum is not solely emitted by the accretion disk. Korista & Goad (2001) first noted that radiation reprocessed by the BLR contributes not only to the formation of emission lines but also to the diffuse continuum emission. This finding was later confirmed by Fausnaugh et al. (2016) and further emphasized by Cackett et al. (2021); Netzer (2022), who demonstrated that the diffuse BLR contribution can significantly dominate the measured continuum time delays. This effect poses challenges for using continuum time delays as a method to determine the Hubble constant, , as originally proposed by Collier et al. (1999). For instance, Cackett et al. (2007) derived an value of km s-1 Mpc-1; however, at the time, not all aspects of the physical mechanisms driving continuum time delays were fully understood.
@@
This paper aims to test a specific model of BLR formation based on radiation pressure acting on dust in disk regions where the disk atmosphere temperature falls below the sublimation threshold. This model, known as the Failed Radiatively Accelerated Dusty Outflow (FRADO), was proposed to explain the vertical rise of material responsible for low-ionization lines, such as H (Czerny & Hryniewicz, 2011; Naddaf et al., 2021; Naddaf & Czerny, 2022). In this study, we investigate the reprocessing of central radiation through a combination of two reprocessors: the Keplerian disk and the BLR, using the BLR response derived from the FRADO model. This approach provides a three-dimensional distribution of BLR clouds, which we then combine with a spectral shape model of the BLR generated using the CLOUDY code (version 23.00; Chatzikos et al. 2023). We fit this model to the observed spectrum and time delays of NGC 5548 and evaluate its potential for determining the distance to the source and the Hubble constant.
@@
The structure of this paper is as follows. In Section 2, we describe the data used for fitting. Section 3 outlines our methodology, which involves generating the response function of the accretion disk using the lamp-post model and incorporating the BLR contribution via FRADO and CLOUDY to estimate the combined time delays. Section 4 presents the results obtained from different models, detailing the simultaneous fitting of the time lag and spectral energy distribution (SED) while assessing the strengths and limitations of each approach. Finally, in Section 5, we estimate the luminosity distance based on our best-fitting model.
@@ 2 Spectral shape and continuum time delay data of NGC 5548
NGC 5548 ( = 0.01651) is one of the most extensively studied AGNs across multiple wavelengths. It was the primary target of the AGN Watch international collaboration (Peterson et al., 2002), which began investigating its emission line variability as early as 1987. Over the years, its optical and UV variability has been the subject of numerous studies (Peterson et al., 1991; Korista et al., 1995; Peterson et al., 1992, 2002, 2004). Observations from both ground-based and space-based telescopes have provided critical insights into accretion disk dynamics and emission line variability. Notably, optical variability in both the continuum and emission lines was first reported by de Vaucouleurs & de Vaucouleurs (1972) in the 1970s, while early UV observations commenced in 1988 with the International Ultraviolet Explorer (IUE; Ulrich & Boisson 1983). More recently, intensive multi-wavelength monitoring campaigns (Edelson et al., 2015; De Rosa et al., 2015; Fausnaugh et al., 2016; Pei et al., 2017; Horne et al., 2021; Lu et al., 2022) have significantly deepened our understanding of the source.
@@
These studies have revealed that the measured continuum time delays do not follow a simple power-law trend. Instead, they show clear evidence of interaction with an extended BLR, where different emission lines originate. Consequently, these delays are often approximated by mean time delays, though some emission line delay profiles exhibit double-peaked features. This complexity opens up new avenues for testing specific models of the BLR structure and response.
@@
In this study, we utilize the broad-band SED and global parameters from Mehdipour et al. (2015), covering the period from 2013 to 2014. The broad-band SED, corrected for internal extinction, starlight contamination, the Balmer continuum, and the Fe II pseudo-continuum, was first presented by Mehdipour et al. (2015). Subsequently, Kubota & Done (2018) fitted this spectrum to an AGN spectral model, revealing that only the outermost part of the accretion flow corresponds to a standard disk. Their results further indicated that the optical/UV emission primarily originates from a region dominated by the warm corona (see their Figure 4).
@@
Therefore, to systematically model the observed inter-band continuum time delays, we propose a step-by-step approach. First, we assume a standard accretion disk model (Model A). Next, we consider an alternative scenario (Model B) based on the spectral decomposition proposed by Kubota & Done (2018), in which optical/UV variability arises solely from the reprocessing of X-rays in the outer cold disk. In this case, the warm corona, with a temperature of approximately K, is too optically thick to absorb and re-emit incident hard X-rays in the UV/X-ray bands, although it still contributes to the UV continuum. Finally, we introduce Model C, a new spectral decomposition based on Figure 5 of Mehdipour et al. (2015). Since the nature of disk reprocessing may strongly depend on the physical properties of the accretion disk, we also conduct a detailed re-analysis of the SED of the source. The disk parameters used for Models A, B, and C are summarized in Table 1.
@@
To test these models, we utilize the time delay data from Fausnaugh et al. (2016), obtained during the STORM campaign from December 2013 to August 2014. This campaign integrated optical light curves data from 16 ground-based observatories, across the filters, as well as the SDSS filters, coupled with ultraviolet data from the Hubble Space Telescope (HST) and Swift instruments. The inter-band time delays were measured relative to the HST light curve at 1367 Å, providing a crucial dataset for evaluating the proposed disk reprocessing scenarios.
@@ 3.1 Hot and warm corona
The hot corona, responsible for hard X-ray emission, was modeled by Mehdipour et al. (2015). In our approach, we adopt the proposed power-law shape and normalization for the observed spectrum when fitting the warm corona. However, we neglect the contribution of hard X-rays to the optical/UV part of the spectrum.
Unlike Mehdipour et al. (2015), we model the warm corona with a different approach. In our framework, the inner and outer radii of the warm corona are treated as parameters. The outer radius of the warm corona coincides with the inner radius of the standard (outer) disk. The Comptonization process in the warm corona is computed using the analytical formulae from Sunyaev & Titarchuk (1980) and is parameterized by the optical depth and electron temperature. Since the warm corona covers the inner disk, we do not assume a single temperature for the soft photons. Instead, we determine the temperature of the underlying cold disk at each radius and adjust it by the Comptonization amplification factor. These calculations are performed iteratively, ensuring that the soft photon temperature spans a continuous range while preserving the total energy from the accretion flow. This methodology aligns more closely with the approach of Kubota & Done (2018). The computational code used for this purpose was originally presented in Czerny et al. (2003).
@@ 3.2 Cold disk reprocessing
The irradiating flux is thermalized only in the outer, cold regions of the standard accretion disk. To further explore this process, we simulate the lamp-post model to generate the disk’s response function, assuming a pulse duration from the corona of 0.05 days. In this simulation, we exclude corrections for general relativity and energy-dependent reflection effects, as introduced by Kammoun et al. (2023, 2024); Papoutsis et al. (2024), to maintain a relatively simple code suitable for data fitting. As a result, the incident radiation from the corona is assumed to be fully absorbed by the disk, leading to a localized increase in temperature.
To accurately model the disk structure, we employ a non-uniform radial grid extending from the innermost stable circular orbit () to the outer disk radius (). The radial grid spacing follows the relation , where is the radial distance from the central source in units of gravitational radius (). For each radial position, the azimuthal angle is sampled using the relation , where ensures sufficient resolution across the entire radial range. The surface element for each coordinate is defined as in units of . To facilitate further calculations, we convert into Cartesian coordinates, assuming a geometrically thin disk with negligible height.
For each coordinate, we compute the total time delay, , as the sum of two components: the delay from the corona to the disk, , and the delay from the disk to the plane intersecting the equatorial plane at , denoted as . The delay is influenced by key parameters, including the corona height (), the inclination angle of the system, and the black hole mass. By default, we assume a cold Keplerian disk extending down to the ISCO at , with an outer radius of . However, we also consider cases with different inner and outer radii, which can be set through the parameter .
For a given delay, we calculate the flux with only a non-irradiated disk and a combination of disk emission and irradiation.
We calculate the blackbody emission both based solely on the disk temperature and from the combined effects of the disk and irradiation. To obtain the final emission, we subtract the contribution from the disk.
@@
Next, we generate the response function by sending a very short light pulse of = 0.05 days duration and normalizing the result by the incident bolometric luminosity. This is expressed by the following equation:
The time delay in the disk has been described previously by Jaiswal et al. (2023). Our adopted method, while simpler than the approaches presented by Kammoun et al. (2021b) and Kammoun et al. (2021a), which include corrections for general relativity and disk albedo, is sufficiently accurate for the current purpose.

### v3

@@ Abstract
Methods. We adopt the standard accretion disk model, with the option of an inner hot flow, and employ the lamp-post model to account for disk irradiation. Additionally, we model the BLR structure based on radiation pressure acting on dust. The model is parameterized by the black hole mass, (which is fixed), the accretion rate, the viewing angle, the height of the lamp-post, the cloud density, and the cloud covering factor. The resulting continuum time delays arise from a combination of disk reprocessing and the reprocessing of a fraction of radiation by the BLR.
Results. Our model reasonably reproduces the observed broad-band continuum, the H time delay, and the continuum inter-band time delays measured during the observational campaign. When the accretion rate is not constrained by the known distance to the source, our approach allows for a direct estimation of the distance. The resulting Hubble constant, = km s-1 Mpc-1, represents a significant improvement over previously reported values derived from continuum time delays in the literature.
@@
The central components of active galactic nuclei (AGNs) include a massive black hole, a compact X-ray emitting region, a relatively cold Keplerian accretion disk, and the broad line region (BLR) (see, e.g., Krolik, 1999a, for a basic compendium). The Keplerian disk serves as the primary source of continuum emission, while the BLR predominantly produces broad emission lines, along with a minor contribution to the continuum. In terms of spatial extent, the innermost regions of the accretion disk span approximately 10100 gravitational radii (), whereas the BLR extends outward at distances of order of 1000 . Beyond the BLR, a dusty molecular torus is located at even larger radii ( to ). Since the early studies of active galaxies by Seyfert (1943) and the discovery of quasars by Schmidt (1963), there has been tremendous progress in our understanding of the structure of AGNs (see e.g. Krolik, 1999b, for a review).
The innermost regions of an AGN are too compact ( a few microarcseconds in the case of BLR) to be spatially resolved directly. However, their structure can be probed using intrinsic flux variability. Among the various wavelengths, X-ray variability is the fastest, originating from the most compact region near the black hole. The reprocessing of this X-ray emission at longer wavelengths, observed through light echo studies, provides insights into the structure and dynamics of the surrounding material, a technique known as reverberation mapping (RM; Blandford & McKee, 1982; Peterson, 1993). RM has been widely applied to AGNs (Sergeev et al., 2005; McHardy et al., 2014; Mudd et al., 2018; Homayouni et al., 2019; Yu et al., 2020; Guo et al., 2022), particularly in studies of continuum variability, which is believed to arise from the irradiated accretion disk.
Although our understanding of AGN structure is still evolving, significant progress has enabled the exploration of AGNs as potential tools for cosmology. Unlike standard candles, AGNs exhibit luminosities that span several orders of magnitude. Nevertheless, the luminosity distance to individual AGNs can be estimated either through geometric methods or by standardizing them in a statistical sense (see e.g. Czerny et al., 2018, for reviews). Early successful approaches relied on the non-linear relation between UV and X-ray luminosities (Risaliti & Lusso, 2015; Lusso et al., 2025), or the correlation between the BLR radius and the source luminosity (Mart´ınez-Aldama et al., 2019; Cao et al., 2022, 2024, 2025). However, these methods require external scaling and, without it, can only constrain the curvature of the Hubble diagram. As such, they are not directly suitable for determining the Hubble constant.
The Hubble constant can instead be derived using strong gravitational lensing techniques (see Suyu et al., 2024, for a recent review), or through a combination of RM of the BLR with angular size measurements obtained from high-resolution interferometric observations (Li et al., 2022, 2025). While both methods are promising, they demand extensive, dedicated observational campaigns and are currently feasible only for a limited number of AGNs, specifically, nearby, bright sources at low redshift where sufficient spatial resolution can be achieved.
The primary goal of this work is to develop a practical method for measuring the Hubble constant using a single mean spectrum of an AGN combined with photometric RM results at multiple wavelengths. High-quality photometric RM measurements are already available for a number of sources (Fausnaugh et al., 2016; Cackett et al., 2018; Edelson et al., 2024; Prince et al., 2025), and a substantial influx of data is expected from large-scale time-domain surveys, such as the Zwicky Transient Facility (ZTF) and the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST).
The idea of using photometric RM to measure the Hubble constant is not new; it was first proposed by Collier et al. (1999), but has yet to produce successful results. The method is based on the standard accretion disk model of Shakura & Sunyaev (1973), which predicts a direct relationship between the time delay at a given wavelength and the monochromatic luminosity: intrinsically brighter sources are physically larger and thus produce longer time delays. More specifically, the model anticipates that the time delays () should scale with wavelength () as . While all coefficients in this relation are determined by the model, the viewing angle () to the source remains a free parameter and must be estimated independently.
While observational results have roughly confirmed this trend, they have not matched the predicted normalization, as observed delays continue to exceed theoretical expectations (e.g. Cackett et al., 2007; Shappee et al., 2014; Kokubo, 2018; Guo et al., 2022; Mandal et al., 2025). This discrepancy, known as the accretion disk size problem, was first recognized in optical microlensing studies of gravitationally lensed quasars (Morgan et al., 2010) and till recently remained an ongoing challenge in our understanding of AGN accretion disk structure.
@@
Recent intensive RM campaigns (Edelson et al., 2015; De Rosa et al., 2015; Fausnaugh et al., 2016; Pei et al., 2017; Horne et al., 2021; Lu et al., 2022; Prince et al., 2025) have revealed that the observed optical continuum is not solely emitted by the accretion disk. Korista & Goad (2001) first noted that radiation reprocessed by the BLR contributes not only to the formation of emission lines but also to the diffuse continuum emission. This finding was later confirmed by Fausnaugh et al. (2016) and further emphasized by Cackett et al. (2021); Netzer (2022), who demonstrated that the diffuse BLR contribution can significantly dominate the measured continuum time delays. This effect poses challenges for using continuum time delays as a method to determine the Hubble constant, , as originally proposed by Collier et al. (1999). For instance, Cackett et al. (2007) derived an value of km s-1 Mpc-1; however, at the time, not all aspects of the physical mechanisms driving continuum time delays were fully understood.
The contribution of the BLR to the continuum time delay implies that, in order to use photometric continuum RM of AGNs for measuring the Hubble constant, we require not only a reliable accretion disk model but also a robust model of the BLR structure. Importantly, such a BLR model must be fully determined by the absolute luminosity of the source. However, fully parametric BLR models (e.g. Baldwin et al., 1995; Li et al., 2013; Pancoast et al., 2014; Grier et al., 2017) may not be adequate in this context, as the ability to recover the absolute luminosity of the source can be compromised unless wavelength-resolved spectroscopic RM data are available. Without such spectral resolution, key constraints on the BLR geometry and its luminosity dependence may be lost, limiting the accuracy of distance measurements based on photometric continuum RM.
@@
We thus propose the use of the Failed Radiatively Accelerated Dusty Outflow (FRADO) model of the BLR (Czerny & Hryniewicz, 2011; Naddaf et al., 2021; Naddaf & Czerny, 2022). The model assumes BLR formation based on radiation pressure acting on dust in disk regions where the disk atmosphere temperature falls below the sublimation threshold. It was proposed to explain the vertical rise of material responsible for low-ionization lines, such as H. In this study, we thus investigate the reprocessing of central radiation through a combination of two reprocessors: the Keplerian disk, which we modify to include the warm corona in the innermost part, and the BLR, using the BLR response derived from the FRADO model. This approach provides a three-dimensional distribution of BLR clouds, which we then combine with a spectral shape model of the BLR generated using the CLOUDY code (version 23.00; Chatzikos et al. 2023a). We fit this model to the observed spectrum and time delays of NGC 5548 and evaluate its potential for determining the distance to the source and the Hubble constant. We test the model for a specific source, and we treat the results as the pilot study: a base for further model developments and the later use for other objects.
NGC 5548 ( = 0.017175), a Seyfert 1 galaxy, is among the most extensively studied AGNs. Its brightness, relatively close distance, and significant variability made it an early candidate for studying the inner, unresolved structure of an active nucleus (e.g., Clavel et al., 1991; Peterson et al., 1991; Krolik et al., 1991; Rokaki et al., 1993). The recent high-quality photometric monitoring (Fausnaugh et al., 2016), broad-band spectral analysis (Mehdipour et al., 2015), measured H time delay from spectroscopic RM (Pei et al., 2017), and wavelength-resolved RM that captured the detailed response of the BLR (Horne et al., 2021) collectively make NGC 5548 an ideal laboratory for testing our methodology.
@@
The structure of this paper is as follows. In Section 2, we describe the data used for fitting. Section 3 outlines our methodology, which involves generating the response function of the accretion disk using the lamp-post model and incorporating the BLR contribution via FRADO and CLOUDY to estimate the combined time delays. Section 4 presents the results obtained from different models, detailing the simultaneous fitting of the time lag and spectral energy distribution (SED) while assessing the strengths and limitations of each approach. Finally, in Section 5, we estimate the luminosity distance based on our best-fitting model. The discussion and conclusions are presented in Sections 6 and 7, respectively.
@@ 2 Spectral shape and continuum time delay data of NGC 5548
NGC 5548 ( = 0.017) is one of the most extensively studied AGNs across multiple wavelengths. It was the primary target of the AGN Watch international collaboration (Peterson et al., 2002), which began investigating its emission line variability as early as 1987. Over the years, its optical and UV variability has been the subject of numerous studies (Peterson et al., 1991; Korista et al., 1995; Peterson et al., 1992, 2002, 2004). Observations from both ground-based and space-based telescopes have provided critical insights into accretion disk dynamics and emission line variability. Notably, optical variability in both the continuum and emission lines was first reported by de Vaucouleurs & de Vaucouleurs (1972) in the 1970s, while early UV observations commenced in 1988 with the International Ultraviolet Explorer (IUE; Ulrich & Boisson 1983). More recently, intensive multi-wavelength monitoring campaigns (Edelson et al., 2015; De Rosa et al., 2015; Fausnaugh et al., 2016; Pei et al., 2017; Horne et al., 2021; Lu et al., 2022) have significantly deepened our understanding of the source.
@@
These studies have revealed that the measured continuum time delays do not follow a simple power-law trend with wavelength. Instead, they show clear evidence of interaction with an extended BLR, where different emission lines originate. Consequently, these delays are often approximated by mean time delays, though some emission line delay profiles exhibit double-peaked features. This complexity opens up new avenues for testing specific models of the BLR structure, as well as for explaining the observed inter-band continuum time delays.
@@
In this study, we utilize the broadband SED and global parameters from Mehdipour et al. (2015), covering the period from 2013 to 2014. The broadband SED, corrected for internal extinction, starlight contamination, the Balmer continuum, and the Fe II pseudo-continuum, was first presented by Mehdipour et al. (2015). Subsequently, Kubota & Done (2018) fitted this spectrum to an AGN spectral model, revealing that only the outermost part of the accretion flow corresponds to a standard disk. Their results further indicated that the optical/UV emission primarily originates from a region dominated by the warm corona (see their Figure 4).
Therefore, to systematically model the observed inter-band continuum time delays, we propose a step-by-step approach. We select three representative setups to illustrate how the spectral fit and time-delay measurements depend on the pre-set parameters and details of the geometry.
First, we assume a standard accretion disk model (Model A). This choice is motivated by the delay fitting for NGC 5548 done by Kammoun et al. (2021b). The source was fitted by Kammoun et al. (2021b) without assuming an inner hot flow, only with the corona height, black hole mass (), spin, accretion rate, and the X-ray luminosity as free parameters. The color correction was set to 2.4. We modify the approach: we fix and the Eddington ratio, but we allow for the inner hot flow and the contribution of the BLR. Our color correction is the same, and no warm corona is present. Since we also fit the spectral shape, we include a starlight component in the model.
@@
Next, we consider an alternative scenario (Model B) based on the spectral decomposition proposed by Kubota & Done (2018). In this case, the warm corona plays an important role in the spectral fit. We adopt the value of the warm corona temperature from the corresponding paper. In order to calculate the time delay caused by reprocessing, we assume that the optical/UV variability arises solely from the reprocessing of X-rays in the outer cold disk. This is motivated by the fact that warm corona with a temperature of approximately K, is too optically thick to absorb and re-emit incident hard X-rays in the UV/X-ray bands, although it still contributes to the UV continuum. The model is supplemented with the BLR contribution to the delays and the spectrum, and starlight is included.
Finally, we introduce Model C, a most general spectral decomposition based on Figure 5 of Mehdipour et al. (2015). Now a number of geometrical and physical parameters are optimized to fit the spectral and time delay data. This includes the warm corona temperature and its optical depth, as well as contribution from BLR and starlight.
The disk parameters used for Models A, B, and C are summarized in Table 1, with more detailed descriptions provided in Section 4.4.
@@
To test these models, we utilize the continuum time delay data from Fausnaugh et al. (2016), obtained during the STORM campaign from December 2013 to August 2014. This campaign integrated optical light curves data from 16 ground-based observatories, across the filters, as well as the SDSS filters, coupled with ultraviolet data from the Hubble Space Telescope (HST) and Swift instruments. The inter-band time delays were measured relative to the HST light curve at 1367 Å, providing a crucial dataset for evaluating the proposed disk reprocessing scenarios.
@@ 3.1 Hot and warm corona
The hot corona, responsible for hard X-ray emission, was modeled by Mehdipour et al. (2015). In our approach, we adopt the proposed power-law shape and normalization for the observed spectrum when fitting the warm corona. We do not model the hot corona; instead, the power-law component is treated as part of the broadband spectrum that irradiates the BLR clouds, as discussed in the next section. It is also included in the fit to the global spectrum, in the X-ray region. However, we neglect the contribution of hard X-rays to the optical/UV part of the spectrum.
Unlike Mehdipour et al. (2015), we model the warm corona with a different approach. In our framework, the inner and outer radii of the warm corona are treated as parameters. The outer radius of the warm corona coincides with the inner radius of the standard (outer) disk. The Comptonization process in the warm corona is computed using the analytical formulae from Sunyaev & Titarchuk (1980) and is parameterized by the optical depth and electron temperature. Since the warm corona covers the inner disk, we do not assume a single temperature for the soft photons. Instead, we determine the temperature of the underlying cold disk at each radius and adjust it by the Comptonization amplification factor. These calculations are performed iteratively, ensuring that the soft photon temperature spans a continuous range while preserving the total energy from the accretion flow. The coronal parameters (optical depth and the temperature) are assumed to be radius-independent as we do not have a firm predictions for these parameters from the warm corona physics. This methodology aligns closely with the approach of Kubota & Done (2018). The computational code used for this purpose was originally presented in Czerny et al. (2003). The warm corona is thus fully modeled, with its spectral shape and normalization determined by global fitted parameters, including the optical depth, temperature, the disk radius at which the warm corona develops, and the accretion rate. This is incorporated in the spectral fitting, but not included separately in BLR reprocessing where the incident radiation is made at the basis of the broad band SED, instead of a separate modeling of the disk and the warm corona. This choice avoids the computational complexity of calculating the reprocessed emission using the CLOUDY code (Chatzikos et al., 2023b) at every step of the simulation. It is stated explicitly later on.
@@ 3.2 Cold disk reprocessing
The irradiating flux is thermalized only in the outer, cold regions of the standard accretion disk. To further explore this process, we simulate the lamp-post model to generate the disk’s response function, assuming a pulse of short duration. In this simulation, we exclude corrections for general relativity (GR) and energy-dependent reflection effects, as introduced by Kammoun et al. (2023, 2024a); Papoutsis et al. (2024), to maintain a relatively simple code suitable for data fitting. As a result, the incident radiation from the corona is assumed to be fully absorbed by the cold disk, leading to a localized increase in temperature. We briefly comment on the implications and limitations of this simplification in Section 6.
To accurately model the disk reprocessing, we employ a non-uniform radial grid extending from the innermost stable circular orbit () to the outer disk radius (), ensuring sufficient resolution at smaller radii. The radial grid spacing follows the relation , where is the radial distance from the central source in units of gravitational radius, . For each radial position, the azimuthal angle is sampled using the relation , where ensures sufficient resolution across the entire radial range. The surface element for each coordinate is defined as in units of . To facilitate further calculations, we convert into Cartesian coordinates, assuming a geometrically thin disk with negligible height.
For each coordinate, we compute the total time delay, , as the sum of two components: the delay from the corona to the disk, , and the delay from the disk to the plane intersecting the equatorial plane at (), denoted as . The delay is influenced by key parameters, including lamp-luminosity (), , accretion rate (), the corona height (), and the inclination angle of the system (). By default, we assume a cold Keplerian disk extending down to the ISCO at , with an outer radius of . However, we also consider cases with different inner and outer radii, which can be set through the parameter .
The non-irradiated disk is characterized by its flux, , and temperature, , as defined in equations 1 and 2, respectively. When the contribution from the corona is included, the resulting total flux and temperature are given by equations 3 and 4. For a specified delay, we use the combined flux from both the disk emission and irradiation, , as described in equation 3. We then convert this total flux into an effective temperature, , according to equation 4. This temperature is subsequently used to compute the blackbody emission.
Additionally, we account for a color correction factor to the disk temperature, which raises the temperature of the disk atmosphere to , where corresponds to no color correction applied, as described below (Shimura & Takahara, 1995):
@@
Next, we generate the response function by sending a very short light pulse of = 0.05 days duration and normalizing the result by the incident bolometric luminosity. The pulse duration is chosen to balance accuracy and temporal resolution: a very short pulse can introduce significant errors in the temperature enhancement, while a very long pulse tends to smear the signal and reduce time resolution, especially at the shortest wavelengths. The response function for a bare disk is given by the following equation:
The time delay in the disk has been described previously by Jaiswal et al. (2023). Our adopted method, while simpler than the approaches presented by Kammoun et al. (2021b) and Kammoun et al. (2021a), which include corrections for GR and disk albedo, is sufficiently accurate for the current purpose of this pilot study.